if a man weight 900N on earth,what is his weight on jupiter where the acceleration due to gravity is 25.9 m/s



1
Expert's answer
2021-02-11T17:09:45-0500

The mass of the man is:


"m = \dfrac{P_e}{g_e}"

where "P_e = 900N" is his weight on Earth, and "g_e = 9.81m\/s^2" is the gravitational acceleration on Earth. The weight of the man on Jupiter is:


"P_j = mg_j = \dfrac{P_eg_j}{g_e}"

where "g_j = 25.9m\/s^2" is the gravitational acceleration on Jupiter. Thus, obtain:


"P_j = \dfrac{900N\cdot 25.9m/s^2}{9.81m/s^2} \approx 2376N"

Answer. 2376 N.